# boxthm]Proposition boxthm]Example boxthm]Corollary boxthm]Lemma boxthm]Definition boxthm]Assumption

### v1

@@ Abstract
Ensuring fairness in matching algorithms is a key challenge in allocating scarce resources and positions. Focusing on Optimal Transport (OT), we introduce a novel notion of group fairness requiring that the probability of matching two individuals from any two given groups in the OT plan satisfies a predefined target. We first propose FairSinkhorn, a modified Sinkhorn algorithm to compute perfectly fair transport plans efficiently. Since exact fairness can significantly degrade matching quality in practice, we then develop two relaxation strategies. The first one involves solving a penalised OT problem, for which we derive novel finite-sample complexity guarantees. This result is of independent interest as it can be generalized to arbitrary convex penalties. Our second strategy leverages bilevel optimization to learn a ground cost that induces a fair OT solution, and we establish a bound guaranteeing that the learned cost yields fair matchings on unseen data. Finally, we present empirical results that illustrate the trade-offs between fairness and performance.
@@ 1 Introduction
Algorithmic matching mechanisms play an increasing role in modern societies, handling the distribution of rare goods by connecting individuals or firms through algorithmic decisions rather than through price-driven markets. Instead of choosing an exchange partner through a decentralized market in which prices make matches, participants are allocated one or multiple partners through a centralized clearing house that takes into account their individual characteristics and preferences. Examples of such mechanisms include online job recommendations, college admissions systems, dating and ride-hailing apps, and kidney allocation circuits.
@@
An increasing concern is the fairness of such mechanisms: since matchings partially determine individual outcomes (for instance, through marriage or schooling decisions), they may result in unfair outcomes. For instance, the growing importance of dating apps in contemporary matrimonial dynamics has raised concerns over their role in fostering social and racial marital homogamy [zheng2018fairness, jia2018online, zhao2024leveraging], a phenomenon often highlighted as a key driver of inequality. Similarly, school admission systems have been criticized for excessively matching students from privileged backgrounds with elite institutions regardless of their academic potential [hiller2012choix, fack2014impact, simioni2022comment]—causing underprivileged students to lose the benefits of elite higher education. More generally,
these concerns arise because matching decisions overly depend on individual characteristics that are highly correlated with attributes defining social groups, such as ethnicity or socioeconomic background.
Such dependence may lead to highly homogeneous pairings, which may conflict with political, philosophical, and legal principles that promote diversity and equality of opportunity. Algorithmic decisions systematically disadvantaging certain groups are often referred to as group fairness issues.
@@
Group fairness has been extensively studied in supervised and unsupervised learning [barocas-hardt-narayanan]. In supervised learning, where an algorithm predicts an unknown outcome $Y$ from features $X$, one example of group fairness is demographic parity [calders2009building], which requires that the prediction $f(X)$ be independent of a protected attribute $S$ (such as gender or age). In practice, this means that predictions for individuals—such as credit scores—should not systematically depend on sensitive attributes. In unsupervised learning, one approach is equality of representation error, which requires that the reconstruction error of a learned latent representation be independent of the sensitive attribute $S$ [samadi2018price]. Other notions of group fairness have also been proposed in both supervised and unsupervised contexts [barocas-hardt-narayanan].
@@
Extending these fairness notions to matching problems is, however, nontrivial, as matching cannot be directly framed as either a supervised or unsupervised learning task. Most prior work has focused on individual fairness (ensuring that similar individuals receive similar matches) or on problem-specific notions, such as guaranteeing a minimal participation level in settings where individuals can opt out of the matching mechanism. We discuss this literature in more details in the related work section.
Contributions.
@@
In this work, we study group fairness in matching using the framework of optimal transport (OT). OT has been a long standing modeling tool in economics and social sciences since the pioneering work of Kantorovitch, eventually earning him the Nobel Memorial Prize in Economic Sciences in 1975 [galichon2018optimal, galichon2021unreasonable]. It has gained traction in the last decade as a way of studying complex matching mechanisms, finding applications in the economics of international trade [galichon2021unreasonable, mastrandrea2025optimal], logistics and urban network studies [leite2022urban], demand estimation and pricing [qu2025sinkhorn], decision theory, as well as labor and migration economics [carlier2023sista, hazard2025whom]. We introduce a novel notion of fairness for optimal transport that allows a central planner—such as a government or regulatory body—to specify the probability that individuals from one subgroup are matched with individuals from another subgroup. This flexible definition can enforce a variety of constraints, including limited homogamy (reducing the likelihood that individuals with similar sensitive attributes are matched), matching quotas (ensuring that $p$ percent of a minority subgroup is matched to another group), or broader goals of social diversity. Specifically, our contributions are as follows.
@@
    We first provide a theoretical framework for fair optimal transport, introducing a formal and novel definition. Building on this framework, we present the FairSinkhorn algorithm for computing optimal transport plans under exact fairness constraints.
@@
    We design two relaxed approaches. The first one penalizes the optimal transport objective with a convex penalty that promotes fair transport plans. This formulation yields a convex optimization problem in which the utility–fairness trade-off can be easily controlled by adjusting the fairness penalty parameter. The second approach consists in learning a cost function such that the resulting transport plan is fair. It has the advantage of enabling to easily match new samples.
@@
    For the first approach, we provide a novel sample complexity bound for penalized entropic optimal transport, which we believe to be of independent interest. For the cost learning approach, we provide a bound on the expected fairness deviation when using finite sample to compute the transport plan with the learned cost function.
@@
    Lastly, we investigate the empirical performance of the proposed approaches through a series of numerical experiments.
@@ 2 Related Work
Fairness in Matching Mechanisms.
Since the pioneering work of gale1962college on the stable marriage problem, a series of works have provided theoretical studies of the fairness in matching mechanisms. A first line of research focuses on the stability of GS-like algorithms under constraints akin to individual fairness, ensuring that similar individuals receive similar outcomes by the decision-making process. For instance, karni2021fairness analyze the Gale-Shapley (GS) algorithm through the lens of preference informed individual fairness. This criterion allows for small deviations from individual fairness, which requires that similar individuals should be treated similarly. They devise a fair variant of the GS algorithm that ensure both approximately fair and stable matchings. Similarly, devic2023fairness analyze merit-based fairness of one-to-one matching mechanisms, a notion close to individual fairness, under partial uncertainty over individual’s merit. A second line of research introduces metrics of fairness closer to group fairness, with a specific focus on kidney paired donation programs, a singular matching problem subject to domain specific constraints — in these programs, donors and patients must be matched by forming cycles due to legal constraints, which leads to a reformulation of the matching problem as a constrained integer programming problem. In their seminal work, ashlagi2014free unveil an unfair market collapse phenomenon in kidney pair donation programs: instead of introducing all patients to the matching system, hospitals internally operate easy matches i.e. do not participate fully in the exchange mechanism. This leads to a loss in transplantation opportunities for highly-sensitized patients that are hard to match. dickerson2014price and following work by zhang2025fairness also consider the case of kidney paired donations programs. Their definition of fairness enforces a minimal amount of program participation for the group of highly-sensitized patients. More recently, lobo2025fair consider a combinatorial multimatching problem under uncertainty: in their setup, the goal is to match agents and resources without exact knowledge of their pairwise valuations. Their definition of fairness relies on the average utility within predefined groups, which is closest to our definition among previous work.
Our contribution differentiates from these works in two aspects. First, we do not consider iterative matching mechanisms such as the GS algorithm: our work focuses on matching through (entropic) optimal transport. Secondly, our definition of fairness is different from previously introduced definitions, and highly modular to accommodate fairness preferences of a central planner. Instead of requiring individual fairness, building on individual rankings over possible assignements or participation constraints, we consider a matching problem between two populations (f.i. students and universities) further structured by a sensitive attribute (f.i. privileged and unprivileged students, and elite and non-elite universities). We enforce matching quotas given by a central planner, that specify for each subgroup the proportion of individuals that should be matched to subgroups in the second population.
Fairness and Optimal Transport.
Closer in spirit to our work is the recent article by nguyen2024towards, which imposes fairness constraints directly on a Wasserstein barycenter by requiring approximate equality of sliced-Wasserstein distances SW [nadjahi2020statistical] to multiple marginals $\mu_{1},\dots,\mu_{K}$:
Their approach, however, does not consider sensitive attributes or demographic groups, focusing instead on balancing distances across marginals. In contrast, we study the problem of computing transport plans that satisfy explicit mass constraints between groups defined by sensitive attributes, a setting that, to the best of our knowledge, has not been previously considered.
Constrained and Statistical Optimal Transport.
3 Formalizing Group Fairness in Optimal Transport
@@ 3.1 Optimal Transport
Entropic Optimal Transport.
We consider the entropic regularized optimal transport problem [peyre2019computational, genevay2019sample, feydy2019interpolating, pooladian2021entropic, keriven2022entropic] between two probability distributions $\mu$ and $\eta$ with ground cost $c$:
$\mathbf{W}_{c,\varepsilon}(\mu,\eta):=\min\limits_{\pi\in\Pi(\mu,\eta)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}\big(\pi|\mu\otimes\eta\big),$
where the minimum is taken over couplings $\Pi(\mu,\eta)$ of $\mu$ and $\eta$ and where $\mathbf{KL}$ denotes the Kullback-Leibler (KL) divergence [CSZISAR_SUBDIFF]. We focus on the entropic formulation because the regularization ensures uniqueness and smoothness of the optimal coupling, which is crucial for efficient computation and gradient-based optimization [peyre2019computational]. With a slight abuse of notation, we will often omit the explicit dependence on $\mu$ and $\eta$ when clear from context, writing, for example, $\mathbf{W}_{c,\varepsilon}$ instead of $\mathbf{W}_{c,\varepsilon}(\mu,\eta)$.
@@
$\mathbf{W}_{\mathbf{C},\varepsilon}:=\min\limits_{\mathbf{\Pi}\in\Pi}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{m}\mathbf{\Pi}_{ij}\mathbf{C}_{ij}+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\kern 5.0pt,$ (1)
where $\Pi:=\big\{\mathbf{\Pi}\in\mathbb{R}_{+}^{n\times m}\,|\,\mathbf{\Pi}\mathbf{1}_{m}=1/n\mathbf{1}_{n},\,\mathbf{\Pi}^{\top}\mathbf{1}_{n}=1/m\mathbf{1}_{m}\big\}$, $\mathbf{KL}\big(\mathbf{\Pi}\big):=\sum_{ij}\mathbf{\Pi}_{ij}\log\mathbf{\Pi}_{ij}$ and $\mathbf{C}\in\mathbb{R}^{n\times m}$ with $\mathbf{C}_{ij}:=c(x_{i},y_{j})$. The KL divergence is strictly convex, and $\Pi$ is a nonempty and bounded set. Therefore, Problem (1) admits a unique minimizer, which we denote by $\mathbf{\Pi}_{\varepsilon}(\mathbf{C})$.
@@ 3.2 Fair Optimal Transport
We consider two distributions $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$ to be matched, where $\mathcal{X}$ and $\mathcal{Y}$ represent feature spaces while $\mathcal{S}$ and $\mathcal{W}$ correspond to sensitive attributes (e.g., gender, ethnicity or age) defining groups of entities in $\mu$ and $\eta$ respectively. The sets $\mathcal{W}$ and $\mathcal{S}$ are assumed to be finite, and are identified, respectively, with $\{1,\dots,K_{w}\}$ and $\{1,\dots,K_{s}\}$. We first make an assumption on the support of the measures, which is needed to leverage several recent results in statistical optimal transport [rigollet2022sample].
@@
The measures $\mu$ and $\eta$ are compactly supported.
We let $p_{s}=\mathbb{P}(S=s)$ and $q_{w}=\mathbb{P}(W=w)$ the probability of each group $i$ in $\mu$ and $j$ in $\eta$, respectively, and denote by $\mathbf{p}$ and $\mathbf{q}$ the resulting distributions over $\mathcal{S}$ and $\mathcal{W}$.
Fairness targets.
Figure 1: Summary of the fairness issues arising in optimal transport when sensitive features are highly correlated with features used for cost computations. On the left of the figure, we show an illustrative setting in which we wish to match two populations (circles and triangles). Group membership is highly correlated with the individual’s location in the $(x,y)$ plane. This results in a block sparse transport plan, displayed on the right.
Optimal transport imposes structure on the coupling of two probability distributions through the cost function, making it more likely to match points whose transportation cost to one another is low. While this inductive bias is often desirable, it can lead to undesirable outcomes when features in $\mathcal{X}$ and $\mathcal{Y}$ are correlated with the sensitive attributes in $\mathcal{S}$ and $\mathcal{W}$, as the resulting coupling may reflect, preserve, or amplify disparities between the corresponding groups, as illustrated in the following example.
@@
Consider a school assignment system, in which students are characterized by their geographic position $X\in\mathbb{R}^{2}$ and a social status $S\in\{\textnormal{high},\textnormal{low}\}$. Similarly, schools are characterized by geographic positions $Y\in\mathbb{R}^{2}$ and a prestige level $W\in\{\textnormal{elite},\textnormal{non-elite}\}$. Suppose that students with $S=\textnormal{high}$ are more likely to live near elite schools, which are often situated in privileged neighborhoods, while students with low social status tend to live near non-elite schools. In this setting, assigning students to schools using optimal transport with Euclidean cost—minimizing the distance between students and schools—will result in highly segregated schools.
@@
To address such issues and define a fair transport plan, we are given a fairness target $\mathbf{F}$, which is a $K_{s}\times K_{w}$ matrix that specifies, for each pair of groups $(s,w)\in\mathcal{S}\times\mathcal{W}$, the desired probability of matching members of group $s$ with group $w$. To be valid, the matrix $\mathbf{F}$ should itself be a coupling between $\mathbf{p}$ and $\mathbf{q}$: it should be non-negative and satisfy the constraints
$\sum\limits_{w=1}^{K_{w}}\mathbf{F}_{sw}=p_{s}\textnormal{ and }\sum\limits_{s=1}^{K_{s}}\mathbf{F}_{sw}=q_{w}$
for all $(s,w)$. We hence write $\Pi(\mathbf{p},\mathbf{q})$ the set of fairness targets.
@@
Consider the segregated schooling system introduced in Example 3.1. Let $p:=\mathbb{P}(S=\textnormal{low})$ and $q:=\mathbb{P}(W=\textnormal{non-elite})$. A newly appointed city administrator wishes to strongly limit social homogamy within public schools, and hence requires $60\%$ of student with low social status to be matched to elite schools. The fairness target can be written as
$\mathbf{F}=\begin{bmatrix}0.4\times p&0.6\times p\\ q-0.4\times p&1-q-0.6\times p&\end{bmatrix}$.
Then, we define a coupling to be $\mathbf{F}$-fair if the amount of mass moved from group $s$ to group $w$ matches the fairness target $\mathbf{F}_{sw}$.
Definition 3.3 ($\mathbf{F}$-Fair coupling).
Let $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$. We define the $\mathbf{F}$-fair transport plan between $\mu$ and $\eta$ as
$\pi^{\star}_{c,\varepsilon,\textnormal{fair}}\in\operatorname*{arg\,min}\limits_{\pi\in\Pi_{\textnormal{fair}}(\mu,\eta)}\int c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}(\pi|\mu\otimes\eta)\kern 5.0pt.$ (2)
where the set of fair couplings $\Pi_{\textnormal{fair}}(\mu,\eta)$ is defined as
$\Pi_{\textnormal{fair}}(\mu,\eta):=\Big\{\pi\in\Pi(\mu,\eta)\mid\forall(s,w)\in\mathcal{S}\times\mathcal{W},\,\pi_{SW}(s,w)=\mathbf{F}_{sw}\Big\}$
where $\pi_{SW}$ is the coupling on $\mathcal{S}\times\mathcal{W}$ that is obtained from $\pi$ by marginalizing over $x$ and $y$, that is
$\pi_{SW}(s,w):$ $=\pi(\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\})=\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi(x,u,y,t)\kern 5.0pt.$
In the finite-sample case, where we have access to two datasets $(\mathbf{x}_{i},\mathbf{s}_{i})_{i=1}^{n}\in\mathcal{X}\times\mathcal{S}$ and $(\mathbf{y}_{i},\mathbf{w}_{i})_{i=1}^{m}\in\mathcal{Y}\times\mathcal{W}$ drawn i.i.d. from $\mu$ and $\eta$ respectively, the fair optimal transport problem writes
$\begin{aligned} &\min\limits_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\\ \textnormal{s.t.}\,\,&\forall(s,w)\in\mathcal{S}\times\mathcal{W},\,\,\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]=\sum\limits_{i|\mathbf{s}_{i}=s}\sum\limits_{j|\mathbf{w}_{i}=w}\mathbf{\Pi}_{ij}=\mathbf{F}_{sw}\kern 5.0pt\end{aligned}$ (3)
where $\mathbf{B}_{sw}:=\big(\mathds{1}_{\mathbf{s}_{i}=s}\mathds{1}_{\mathbf{w}_{j}=w}\big)_{i,j}\in\{0,1\}^{n\times m}$. This definition may also be extended to include individual specific weights. The following proposition ensures that this problem is well defined.

### v2

@@ Abstract
Ensuring fairness in matching algorithms is a key challenge in allocating scarce resources and positions. Focusing on Optimal Transport (OT), we introduce a novel notion of group fairness requiring that the probability of matching two individuals from any two given groups in the OT plan satisfies a predefined target. We first propose a modified Sinkhorn algorithm to compute perfectly fair transport plans efficiently. Since exact fairness can significantly degrade matching quality in practice, we then develop two relaxation strategies. The first one involves solving a penalized OT problem, for which we derive novel finite-sample complexity guarantees. Our second strategy leverages bilevel optimization to learn a ground cost that induces a fair OT solution, and we establish a bound on the deviation of fairness when matching unseen data. Finally, we present empirical results illustrating the performance of our approaches and the trade-off between fairness and transport cost.
@@ 1 Introduction
Algorithmic matching mechanisms play an increasing role in modern societies, handling the distribution of rare goods by centrally connecting individuals or firms based on their characteristics and preferences rather than through price-driven markets. Examples of such mechanisms include online job recommendations, college admissions systems, dating and ride-hailing apps, and kidney allocation circuits.
@@
An increasing concern is the fairness of such mechanisms: since matchings partially determine individual outcomes (for instance, through marriage or schooling decisions), they may result in unfair outcomes. For instance, the growing importance of dating apps in contemporary matrimonial dynamics has raised concerns over their role in fostering social and racial marital homogamy [zheng2018fairness, jia2018online, zhao2024leveraging], a phenomenon often highlighted as a key driver of inequality. Similarly, school admission systems have been criticized for excessively matching students from privileged backgrounds with elite institutions regardless of their academic potential [hiller2012choix, fack2014impact, simioni2022comment], causing underprivileged students to lose the benefits of elite higher education. More generally, these concerns arise because matching decisions overly depend on individual characteristics that are highly correlated with attributes defining social groups, such as ethnicity or socioeconomic background. Such dependence may lead to highly homogeneous pairings, which may conflict with political, philosophical, and legal principles that promote diversity and equality of opportunity. Algorithmic decisions systematically disadvantaging certain groups are often referred to as group fairness issues.
@@
Group fairness has been extensively studied in supervised and unsupervised learning [barocas-hardt-narayanan]. In supervised learning, where an algorithm predicts an unknown outcome $Y$ from features $X$, one example of group fairness is demographic parity [calders2009building], which requires that the prediction $f(X)$ be independent of a protected attribute $S$ (such as gender or age). In unsupervised learning, one approach is equality of representation error, which requires that the reconstruction error of a learned latent representation be independent of the sensitive attribute $S$ [samadi2018price]. Other notions of group fairness have also been proposed in both supervised and unsupervised contexts [barocas-hardt-narayanan].
@@
Extending these fairness notions to matching problems is, however, nontrivial, as matching cannot be directly framed as either a supervised or unsupervised learning task. Most prior work has focused on individual fairness, ensuring that similar individuals receive similar matches [devic2023fairness], or on problem-specific notions, such as guaranteeing a minimal participation level in settings where individuals can opt out [ashlagi2014free]. We discuss this literature in more details in the related work section.
Figure 1: Illustration of optimal transport under group fairness constraints. Left: Due to the correlation between group attributes ($S$ and $W$) and features, vanilla optimal transport produces transport plans that concentrate most of the mass between matching groups. Center: Our novel notion of fairness is based on fairness targets, defined as the desired group-level matching probabilities specified by a central planner. Right: We propose two approaches for solving this problem: a fairness-penalized optimal transport method, and a cost-learning approach that transforms the input space so that the resulting optimal transport plan aligns as closely as possible with the fairness targets.
@@
Contributions. In this work, we study group fairness in matching using the framework of optimal transport (OT). Since the pioneering work of Kantorovitch, OT has been a long standing modeling tool in economics and social sciences [galichon2018optimal] and has gained traction in the last decade as a way of studying complex matching mechanisms in economics [galichon2021unreasonable, mastrandrea2025optimal], logistics and urban network studies [leite2022urban], demand estimation and pricing [qu2025sinkhorn], decision theory, as well as labor and migration economics [carlier2023sista, hazard2025whom].
We introduce a novel notion of fairness for optimal transport that allows a central planner—such as a government or regulatory body—to specify the probability that individuals from one group are matched with individuals from another group. This flexible definition can enforce a variety of constraints, including limited homogamy (reducing the likelihood that individuals with the same sensitive attributes are matched), matching quotas (ensuring that a certain percentage of a minority group receives desirable opportunities), or broader goals of social diversity. We summarize our contributions below, with an overview provided in Figure 1.
@@
    We introduce a novel fair optimal transport problem, and present the FairSinkhorn algorithm for computing OT plans under exact fairness constraints.
@@
    We propose two relaxed approaches. The first adds a penalty to the optimal transport objective to encourage fair plans, yielding a convex optimization problem where the cost–fairness trade-off is easily controlled via the penalty parameter. The second learns a cost function such that the resulting transport plan is fair, allowing straightforward matching of new samples.
@@
    For the first approach, we provide a novel sample complexity bound for penalized entropic optimal transport, which we believe to be of independent interest. For the cost learning approach, we bound the expected fairness deviation between finite samples and the distribution, holding for any cost in a given parameterized family.
@@
    We investigate the empirical performance of the proposed approaches through numerical experiments.
@@ 2 Related Work
Fairness in matching algorithms. Fairness has been studied across a variety of matching problems. A first line of work investigates individual-level fairness in Gale–Shapley–type matching mechanisms [gale1962college], requiring that similar individuals receive similar outcomes [karni2021fairness, devic2023fairness]. A second line of work focuses on group-level fairness, primarily in the specific case of kidney paired donation (KPD) programs [ashlagi2014free, dickerson2014price, zhang2025fairness, lobo2025fair]. Our work departs from this literature in two key ways. First, we do not consider iterative agent-level mechanisms (in which agents are sequentially paired until the match is stable) but instead model matching through (entropic) optimal transport. Second, rather than individual rankings, we assume that both populations are structured by sensitive attributes and impose explicit inter-group mass constraints specifying how much probability mass should flow between subgroups. Another related line of work studies fairness-constrained bipartite matching on graphs $\big(\mathcal{U}\cup\mathcal{V},\mathcal{E}\big)$, where edges encode feasible matches [godsil1981matchings, zdeborova2006number, noiry2021online]. Recent work in this setting includes [castera2024price], who partition one side of the graph ($\mathcal{U}$ or $\mathcal{V})$ into groups and limit the number of vertices that can be matched from each group, and [panda2024individual], who further impose individual fairness constraints. Our work is fundamentally different from theirs since $(i)$ individuals are represented by features rather than graph positions; $(ii)$ we allow mass splitting rather than integral matchings; and $(iii)$ our formulation is symmetric in the two populations.
Fairness and optimal transport. A growing literature uses optimal transport to obtain or characterize fair learning algorithms [gordaliza2019obtaining, gouic2020projection, chzhen2020fair, DBLP:conf/aaai/ChiappaJSPJA20, DBLP:conf/nips/BuylB22, hu2023fairness, DBLP:conf/icml/XianY023, xiong2024fair, divol2024demographic]. A central insight is that fair prediction can often be formulated as a Wasserstein barycenter or projection problem [gouic2020projection, chzhen2020fair, divol2024demographic]. In this line of work, OT is a tool to enforce fairness in downstream prediction tasks. Our perspective is orthogonal: we study the fairness of the transport plan itself. Closest to our problem is nguyen2024towards, who impose fairness on a Wasserstein barycenter problem by enforcing approximate equality of its sliced-Wasserstein distances to the input measures. In contrast, we study the problem of computing transport plans that satisfy explicit mass constraints between groups defined by sensitive attributes, a setting that, to the best of our knowledge, has not been previously considered.
Constrained and statistical optimal transport. Our approach builds on constrained optimal transport, which enforces structural properties of transport plans [courty2016optimal, blondel2018smooth, paty2019subspace, liu2022sparsity], including bounds on transported mass [korman2015optimal]. We build on and extend the theory of statistical and penalized optimal transport [rakotomamonjy2015generalized, genevay2019sample, rigollet2022sample] to derive finite-sample guarantees and to control deviations from the prescribed fairness constraints. We provide an extended related work review in Appendix A.
3 Formalizing Group Fairness in OT
@@ 3.1 Optimal Transport
We consider the entropic optimal transport problem [peyre2019computational, genevay2019sample, feydy2019interpolating, pooladian2021entropic, keriven2022entropic] between two probability distributions $\mu$ and $\eta$ with ground cost $c$:
$\min\limits_{\pi\in\Pi(\mu,\eta)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}\big(\pi|\mu\otimes\eta\big),$
where the minimum is taken over couplings $\Pi(\mu,\eta)$ of $\mu$ and $\eta$ and where $\mathbf{KL}$ denotes the Kullback-Leibler (KL) divergence [CSZISAR_SUBDIFF]. We focus on the entropic formulation as regularization ensures uniqueness and smoothness of the optimal coupling, which is crucial for efficient computation and gradient-based optimization [peyre2019computational].
@@
$\min\limits_{\mathbf{\Pi}\in\Pi}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbf{\Pi}_{ij}\mathbf{C}_{ij}+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\kern 5.0pt,$ (1)
where $\Pi:=\big\{\mathbf{\Pi}\in\mathbb{R}_{+}^{n\times m}\,|\,\mathbf{\Pi}\mathbf{1}_{m}=1/n\mathbf{1}_{n},\,\mathbf{\Pi}^{\top}\mathbf{1}_{n}=1/m\mathbf{1}_{m}\big\}$, $\mathbf{KL}\big(\mathbf{\Pi}\big):=\sum_{ij}\mathbf{\Pi}_{ij}\log\mathbf{\Pi}_{ij}$ and $\mathbf{C}\in\mathbb{R}^{n\times m}$ with $\mathbf{C}_{ij}:=c(x_{i},y_{j})$. The KL divergence is strictly convex, and $\Pi$ is a nonempty and bounded set. Therefore, (1) admits a unique minimizer, which we denote by $\mathbf{\Pi}_{\varepsilon}(\mathbf{C})$.
@@ 3.2 Fair Optimal Transport
We consider two distributions $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$ to be matched, where $\mathcal{X}$ and $\mathcal{Y}$ represent feature spaces while $\mathcal{S}$ and $\mathcal{W}$ correspond to sensitive attributes (e.g., gender, ethnicity or age) defining groups of entities in $\mu$ and $\eta$ respectively. The sets $\mathcal{W}$ and $\mathcal{S}$ are assumed to be finite, and are identified, respectively, with $\{1,\dots,K_{w}\}$ and $\{1,\dots,K_{s}\}$. To obtain theoretical results on sample complexity, we will need an assumption on the support of the measures, which is standard in statistical optimal transport [rigollet2022sample].
@@
$\mu$ and $\eta$ are compactly supported.
We denote by $p_{s}=\mathbb{P}(S=s)$ and $q_{w}=\mathbb{P}(W=w)$ the probability of each group $i$ in $\mu$ and $j$ in $\eta$, respectively, and by $\mathbf{p}$ and $\mathbf{q}$ the resulting distributions over $\mathcal{S}$ and $\mathcal{W}$.
Fairness targets. Optimal transport imposes structure on the coupling of two probability distributions through the cost function, making it more likely to match points whose transportation cost to one another is low. While this inductive bias is often desirable, it can lead to undesirable outcomes when features in $\mathcal{X}$ and $\mathcal{Y}$ are correlated with the sensitive attributes in $\mathcal{S}$ and $\mathcal{W}$, as the resulting coupling may reflect, preserve, or amplify disparities between the corresponding groups, as illustrated in the following example.
@@
Consider a school assignment system, in which students have a location $X\in\mathbb{R}^{2}$ and social status $S\in\{\textnormal{high},\textnormal{low}\}$, and schools have a location $Y\in\mathbb{R}^{2}$ and prestige $W\in\{\textnormal{elite},\textnormal{non-elite}\}$. If high-status students tend to live near elite schools and low-status students near non-elite schools, then optimal transport with Euclidean cost—minimizing the distance between students and schools—will produce highly segregated schools.
@@
To address such issues and define fair transport plans, we are given a fairness target $\mathbf{F}$, which is a $K_{s}\times K_{w}$ matrix that specifies, for each pair of groups $(s,w)\in\mathcal{S}\times\mathcal{W}$, the desired probability of matching members of group $s$ with group $w$. To be valid, the matrix $\mathbf{F}$ should itself be a coupling between $\mathbf{p}$ and $\mathbf{q}$: it should be non-negative and satisfy the following constraints for all $(s,w)$:
$\textstyle\sum{w=1}^{K_{w}}\mathbf{F}_{sw}=p_{s}\textnormal{ and }\sum{s=1}^{K_{s}}\mathbf{F}_{sw}=q_{w}\kern 5.0pt.$
We hence write $\Pi(\mathbf{p},\mathbf{q})$ the set of fairness targets.
@@
Consider the segregated schooling system of Example 3.1. Let $p:=\mathbb{P}(S=\textnormal{low})$ and $q:=\mathbb{P}(W=\textnormal{non-elite})$. A newly appointed city administrator wishes to limit social homogamy within public schools, and hence requires $60\%$ of student with low social status to be matched to elite schools. The fairness target can be written as
$\mathbf{F}=\begin{bmatrix}0.4\times p&0.6\times p\\ q-0.4\times p&1-q-0.6\times p\end{bmatrix}$.
Fair optimal transport. We say that a coupling is $\mathbf{F}$-fair if the amount of mass transported from group $s$ to group $w$ equals the prescribed fairness target $\mathbf{F}_{sw}$. An $\mathbf{F}$-fair optimal transport plan is then an $\mathbf{F}$-fair coupling that minimizes the transport cost among all such couplings.
Definition 3.3 ($\mathbf{F}$-Fair Optimal Transport).
Let $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$. An $\mathbf{F}$-fair optimal transport plans between $\mu$ and $\eta$ is defined as
$\pi^{\star}_{c,\varepsilon,\mathbf{F}}\!\in\!\!\operatorname*{arg\,min}\limits_{\pi\in\Pi_{\mathbf{F}}(\mu,\eta)}\int\!\!c(x,y)\,d\pi(x,y)\!+\!\varepsilon\mathbf{KL}(\pi|\mu\otimes\eta)$ (2)
where the set of $\mathbf{F}$-fair couplings $\Pi_{\mathbf{F}}(\mu,\eta)$ is given by
$\Pi_{\mathbf{F}}(\mu,\eta):=\big\{\pi\in\Pi(\mu,\eta)\mid\forall(s,w),\,\pi_{SW}(s,w)=\mathbf{F}_{sw}\big\}$
with $\pi_{SW}$ denoting the induced coupling on $\mathcal{S}\times\mathcal{W}$ obtained from $\pi$ by marginalizing over $x$ and $y$, that is
$\pi_{SW}(s,w):$ $=\pi(\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\})\kern 5.0pt.$
In the finite-sample case, where we have access to two datasets $(\mathbf{x}_{i},\mathbf{s}_{i})_{i=1}^{n}\in\mathcal{X}\times\mathcal{S}$ and $(\mathbf{y}_{i},\mathbf{w}_{i})_{i=1}^{m}\in\mathcal{Y}\times\mathcal{W}$ drawn i.i.d. from $\mu$ and $\eta$ respectively, the fair optimal transport problem (2) writes
$\begin{aligned} &\min\limits_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\\ \textnormal{s.t.}\,&\forall(s,w),\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]\!=\!\!\sum\limits_{i|\mathbf{s}_{i}=s}\sum\limits_{j|\mathbf{w}_{i}=w}\mathbf{\Pi}_{ij}=\mathbf{F}_{sw}\kern 5.0pt\end{aligned}$ (3)
where $\mathbf{B}_{sw}:=\big(\mathds{1}_{\mathbf{s}_{i}=s}\mathds{1}_{\mathbf{w}_{j}=w}\big)_{i,j}\in\{0,1\}^{n\times m}$.
The following proposition ensures that the fair OT problem is well defined.